OpenStudy (anonymous):

If y = 7x – 5, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? (5 points) {(2, 7), (1, 6), (3, 13)} {(0, –5), (2, 9), (–3, –26)} {(–5, 0), (9, 2), (–26, –3)} {(1, 3), (6, 18), (8, 15)}

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@marshallinwashington

OpenStudy (anonymous):

Did you plug in the values to see if any of the answers work?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

@YungXxGodXx

OpenStudy (anonymous):

I'll show you by checking the first answer. The first ordered pair is (2,7). So x= 2 and y=7 Plug into the equation. 7=7(2)-5 7=14-5 7=9 This is not true, so this can't be an answer. The first set is not an answer. Now check one of the ordered pairs in answer number two. So what you get.

OpenStudy (anonymous):

I think the answer is either b or c

OpenStudy (anonymous):

Will let's check b and c The first order pair in b is (0,-5) So x=0 and y=-5 Plug into the equation. -5=0(7)-5 -5=0-5 -5=-5 So this is looking good. Check the other two order pairs to make sure.

OpenStudy (anonymous):

its B!

OpenStudy (anonymous):

Let's check c The first order pair is (1,3) So x=1 and y=3 Plug into the equation. 3=1(7)-5 3=7-5 3=2 This is not true. So if any of the three ordered pairs in this set are not true, this means this set is not a correct answer. So we have eliminated answer a and c. Answer b looks good but we need to check all three ordered pairs.

OpenStudy (anonymous):

thanks!!!!!!!!!!

OpenStudy (anonymous):

You have the correct answer now. Hope this helped you see what you needed to do. yw.